the condition that roots of px2-px+q=0 are in ratio p:q is

Grade:11

1 Answers

abhishek upadhyay
12 years ago

let one be a & other is b then

a= bp/q

a+b = 1

or bp/q +b =1

or b=q/(p+q) .................... (i)     

& ab=q/p

or b2=q2/p2   ....................(ii)                        now equate (i)&(ii)